OpenStudy (anonymous):

I really need help describing how to transform the graph of f(x)=the cubed root of x+1 into g(x)=-2* the cubed root of x-3

OpenStudy (anonymous):

Do you meen \[f(x)=\sqrt{x+1} \rightarrow g(x)=-2\sqrt{ x-3}\]?

OpenStudy (anonymous):

\[G(\sqrt{x+1} ) =-2\sqrt{ x-3} \] if \[ z =\sqrt{x-1}; \] \[ -2\sqrt{(z^2 - 4)} = -2\sqrt{x-3}\]

OpenStudy (anonymous):

is that what you are looking for?